OpenStudy (anonymous):

find the exact value of cos(2 arctan 4/3) without a calculator? please answer with details and explanations if you can.. i'm really confused and there's a test tomorrow. :(

OpenStudy (anonymous):

\[\cos(2\theta)=2\cos^2(\theta)-1\] is a start

OpenStudy (anonymous):

in this case \(\theta)=\arctan(\frac{4}{3})\)

OpenStudy (anonymous):

so your job is first to find \(\cos(\arctan(\frac{4}{3}))\) do you know how to do that?

OpenStudy (anonymous):

it is the same as the question " if the tangent of \(\theta\) is \(\frac{4}{3}\) find the cosine of \(\theta\) "

OpenStudy (anonymous):

any ideas?

OpenStudy (anonymous):

to find cos (2 archtan 4/3) i draw tan 4/3 and then find the cos on the triangle right? so cos would be 3/5?

OpenStudy (anonymous):

yes!! exactly

OpenStudy (anonymous):

and since \(\cos(\theta)=\frac{3}{5}\) then you know \[\cos(2\theta)=2\left(\frac{3}{5}\right)^2-1\]

OpenStudy (anonymous):

or maybe i should say " since you know \(\cos(\tan^{-1}(\frac{3}{4}))=\frac{3}{5}\) you know \[\cos(2\tan^{-1}(\frac{3}{5}))=2\left(\frac{3}{5}\right)^2-1\]

OpenStudy (anonymous):

okay thanks! but does it matter which cos(2θ) formula i use? any of the cos(2θ) formulas would work for this question right?

OpenStudy (anonymous):

typo there, i meant \[\cos(2\tan^{-1}(\frac{3}{4}))=2\left(\frac{3}{5}\right)^2-1\]

OpenStudy (anonymous):

well lets think for a second the formula \[\cos(2\theta)=2\cos^2(\theta)-1\] contains only cosine if you use the one with sine, you will have to find the sine

OpenStudy (anonymous):

so the answer is "yes" you can use whichever one you like, but you want to make life easy, not hard

OpenStudy (anonymous):

oh! that makes a lot of sense :) so technically, i could use any (2θ) formula as long as i find the necessary trig function (sin/cos/tan) for it?

OpenStudy (anonymous):

okay thank you so much!! :D

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

@satellite73 i have a small question.. if you're not too busy, could you answer this really quick please? -- tan(2 arcsin (-3/5)) -- would i solve this the same way as my original question? like find tan(2 arcsin (-3/5)) first, then use tan(2θ) to solve it?

OpenStudy (anonymous):

nvm, i got it! :D